# Unmasking Illusions: Understanding Human Perception of Audiovisual Deepfakes

### v1

@@ III-C1 Data Selection
III-D Web Platform
The experiment was conducted using our gamification-based web platform. We leveraged the popular open-source web server nginx to engage participants in the task of identifying the authenticity of videos they watched. For backend development, we used the high-performance framework PHP Yii2 and used PHP 7.4 as the developing language. MongoDB, known for its flexibility and performance, was used as a NoSQL database, with MongoDB Compass providing an intuitive graphical user interface (GUI). For efficient code editing, we used Sublime Text.
@@ III-E Metrics
[…]
@@
III-F Participant Motivation
Participant motivation plays a crucial role in maintaining engagement and completion rates in subjective studies, especially when the tasks are extensive or repetitive. The risk of the task becoming boring after a few rounds poses a significant challenge that can strongly influence the design process of the study. To alleviate this problem, three approaches were employed.
First, we offered a monetary reward of \$10 upon completion of the study. This approach not only serves as a powerful incentive for participants to participate, but also contributes to a more effective and rewarding research experience overall. Second, we divided the task into two phases. This division allows participants to experience a sense of accomplishment after completing the first phase. By providing this tangible sense of accomplishment at the midpoint, participants are empowered to complete the second phase with renewed energy and focus. Third, we exposed participants to the correct label after each round. This immediate feedback keeps their interest and helps them track their performance. These strategies maintained engagement and ultimately increased their motivation to complete the study.
@@ III-G AI Models
We evaluated five SOTA audiovisual deepfake detection models. These models were trained on the training set of the FakeAVCeleb dataset and tested on the same test set used in the subjective evaluation. This comparison provides scientific insights into human and machine behavior when detecting audiovisual deepfakes.
@@ IV-A Overall Detection Accuracy across Participants
Fig. 6 shows the Phase I (40 rounds) and Phase II (40 rounds) deepfake detection accuracy of individual participants. From Phase I to Phase II, the average accuracy across all participants increased from 63.30% to 67.98%, resulting in an overall accuracy of 65.64%.
Human performance at detecting audiovisual deepfakes is marginally better than random chance. This paper demonstrates that deepfakes have the ability to deceive the majority of the public. This result is consistent with previous research [11], [16] and [21].
Most participants performed better in Phase II than in Phase I. The paired samples t-test analysis showed that the average accuracy improvement from Phase I to Phase II was significant with a p-value less than 0.0001. The results show that the training is effective and that increased exposure and familiarity help them learn and adapt to the stimuli. Over time, their anxiety may reduce and attention and concentration may increase. Furthermore, after each round, participants were exposed to the correct label of the video. This immediate feedback on their performance reflected their correctness and errors, the former increasing their confidence in the detection, and the latter enhancing their ability to accurately distinguish real content from synthetic content.
@@
Interestingly, we also observed a consistent pattern of authenticity bias, with people being more biased towards detecting videos as real, even though they were well-informed that half of the videos were deepfake. This cognitive bias is evident in Fig. 8, where 59% of videos were perceived as real and 41% were identified as fake. Our findings suggest that people follow the principle of “seeing is believing” and tend to believe visual information, leading to interpreting a video as authentic even if it is not. This bias can have a significant impact on decision-making, as people may rely more on intuition than objective evidence, which can lead to them being deceived by fabricated content.
Furthermore, when participants were prompted to indicate the manipulated modality after identifying a video as fake, they tended to believe that both modalities (audio and video) were manipulated, or that there was only visual manipulation. This tendency may stem from visual primacy or cognitive load. Humans are highly visual creatures, and our brains prioritize visual stimuli over auditory signals. One potential factor may be sensory discrepancy; in some cases, deepfake videos exhibit more pronounced visual discrepancies than auditory ones. Furthermore, processing audio and visual information simultaneously may cause cognitive challenges; therefore, participants found it easier to focus on visual cues, encouraging them to prioritize visual information when detecting manipulations in videos. As shown in Fig. 8, of the 41% of “fake” responses, 16% believed that there were two manipulation modes, 15% believed that there was only visual manipulation, and only 10% believed that there was only auditory manipulation. However, each manipulation accounts for about 17% of all videos, and the numbers are balanced.
@@ IV-B Detection Accuracy under Different Participant Groupings
Notably, participants showed improved performance in phase II regardless of grouping, as shown in Fig. 9. The reasons have been discussed in the previous section.
@@
Next, we analyze the correlation between accuracy and confidence, as this insight is valuable for assessing participants’ degree of metacognitive awareness. Therefore, to estimate whether participants correctly perceived their ability to detect audiovisual deepfakes, we compared their actual detection accuracy with self-reported confidence across all videos. Surprisingly, we found that participant’s self-reported confidence in their answers did not align with their accuracy in detecting audiovisual deepfake videos. Despite expressing high confidence, their actual accuracy was low. As shown in Fig. 11, most participants reported average confidence that was higher than their average accuracy. The average confidence level across all participants was 77.60%, while the average accuracy was 65.64%. This difference means that humans’ overconfidence in deepfake detection will cause deepfakes to bring greater harm to human society.
@@ IV-E Performance of AI Models
Model
Accuracy (%)
LipForensics [41] 92.50
AV-Lip-Sync [8] 87.50
AV-Lip-Sync+ [6] 97.50
CNN-Ensemble [7] 97.50
AVTENet [9] 97.50
@@
Fig. 13 compares the performance of humans and 5 AI models in detecting audiovisual deepfakes. Human performance was well below expectations and marginally above the chance level. However, all AI models performed significantly better than human participants. Arguably, AI models have demonstrated an impressive ability to detect subtle manipulations in audiovisual content that might escape human perception. AI models excel humans at integrating information from multiple modalities, and combining audio and visual modalities in deepfake detection contributes significantly to the effectiveness of AI models.
@@ VI Conclusions and Future Work
In this paper, we present the results of a subjective study aimed at analyzing human behavior in detecting audiovisual deepfake videos and comparing human performance with that of five AI models. We tested 110 human subjects and five AI models using the same set of 40 manually selected videos (20 real and 20 fake). The findings showed that subjects generally found detecting deepfakes to be challenging. They performed marginally better than random chance, even when explicitly informed of the presence of deepfakes. The concern is that deepfake technology has reached a level of reality that could confuse much of the public, especially social media users, while people are overconfident in their ability to determine the authenticity of a video. However, over time, participants seemed to refine their perception and develop a more accurate understanding of their detection capabilities. Additionally, we found that familiarizing participants with the threats posed by deepfakes had no significant impact on their ability to detect deepfakes. Participants’ age, gender, and whether they were native speakers of English had an impact on their detection performance, but not their IT skill level. On the other hand, AI models outperformed human participants at detecting forgeries in videos. Furthermore, deepfake detection methods that integrate multiple modalities are more effective than unimodal methods. In order to reduce the harm of deepfakes to human society, it is crucial to develop deepfake detection technology.
To further expand and refine our understanding of human perception of audiovisual deepfakes, we can implement a test design that measures subjective performance in a more discreet and unbiased way. For example, the applicability of our findings can be generalized by including a broader and more diverse population. Detailed analysis of the auditory and visual components of audiovisual signals will reveal how individuals perceive, interpret and process audio and visual cues in multimodal deepfake videos. Furthermore, exploring the impact of the popularity of people in videos (e.g., celebrities or politicians) may provide valuable insights into enhancing human detection capabilities and understanding how prior knowledge affects the perception of corresponding deepfake content.

### v2

@@ III-E Metrics
Accuracy=Number of Correctly Identified VideosTotal Number of $\text{Accuracy}=\frac{\text{Number of Correctly Identified Videos}}{\text{Total Number of Videos}}\times 100\%.$ (1)
@@
III-F Gamification Concepts
Participant motivation plays a crucial role in maintaining engagement and completion rates in subjective studies, especially when the tasks are extensive or repetitive. The risk of the task becoming boring after a few rounds poses a significant challenge that can strongly influence the design process of the study. We used several gamification concepts to keep participants motivated and engaged. Specifically, four approaches were employed.
  • Incentive (Monetary Reward): We offered a monetary reward or incentive of \$10 upon completion of the study. This approach not only serves as a powerful incentive for participants to participate but also contributes to a more effective and rewarding research experience overall.
  • Points and Scoring: Participants received visual feedback showing the number of rounds completed and points earned through the progress bar and correct identification, thus encouraging continued engagement and motivation. In addition, we divided the task into two phases. This division allows participants to experience a sense of accomplishment after completing the first phase. By providing this tangible sense of accomplishment at the midpoint, participants are empowered to complete the second phase with renewed energy and focus.
  • Progress Tracking: Participants could track their progress throughout the study and maintain their attention and engagement through visual feedback on task completion.
  • Feedback Mechanism: We exposed participants to the correct label after each round. This immediate feedback keeps their interest and helps them track their performance. These strategies maintained engagement and ultimately increased their motivation to complete the study.
@@ III-G AI Models
We evaluated five SOTA deepfake detection models. The selection of these five AI deepfake detection methods for this study was based on their status as recent state-of-the-art models and the availability of their test results on the dataset we utilized. These models were trained on the training set of the FakeAVCeleb dataset and tested on the same test set used in the subjective evaluation. This comparison provides scientific insights into human and machine behavior when detecting audiovisual deepfakes.
@@ IV-A Overall Detection Accuracy across Participants
Fig. 7 shows the Phase I (40 rounds) and Phase II (40 rounds) deepfake detection accuracy of individual participants. From Phase I to Phase II, the average accuracy across all participants increased from 63.30% to 67.98%, resulting in an overall accuracy of 65.64%. Most participants performed better in Phase II than in Phase I. Human performance at detecting audiovisual deepfakes is marginally better than random chance. This study demonstrates that deepfakes have the ability to deceive the majority of the public. This result is consistent with previous research [33, 39], [45], [47].
@@
We performed a two-tailed t-test to see if the difference in average accuracy between Phase I and Phase II was significant. The significance level ($\alpha$) was set at $0.05$. The $t$-value is $-5.61$, and the corresponding $p$-value is $1.52\times 10^{-7}$, which indicates a statistically significant difference in accuracy between the two phases. We hypothesize that people improve performance naturally through practice, familiarity, feedback, pattern exposure, or specific audiovisual cues. To prove this, we also performed a one-tailed t-test, as we hypothesized that the average accuracy of Phase II should improve compared to Phase I. The $t$-value remains $-5.61$, and the $p$-value is $7.62\times 10^{-8}$. This extremely small p-value suggests support for the conclusion that the improvement from Phase I to Phase II is meaningful and statistically significant.
The results show that training is effective and that increased exposure and familiarity help participants learn and adapt to the stimuli. Over time, their anxiety may reduce and attention and concentration may increase. Furthermore, after each round, the participants were exposed to the correct video label. This immediate feedback on their performance reflected their correctness and errors, the former increasing their confidence in the detection, and the latter enhancing their ability to accurately distinguish real content from synthetic content.
@@
Interestingly, we also observed a consistent pattern of authenticity bias, with people being more biased towards detecting videos as real, even though they were well-informed that half of the videos were deepfake. This cognitive bias is evident in Fig. 9, where 59% of videos were perceived as real and 41% were identified as fake. Our findings suggest that people follow the principle of “seeing is believing” and tend to believe visual information, leading to interpreting a video as authentic even if it is not. This bias can have a significant impact on decision-making, as people may rely more on intuition than objective evidence, which can lead to being deceived by fabricated content.
Furthermore, when participants were prompted to indicate the manipulated modality after identifying a video as fake, they tended to believe that both modalities (audio and video) were manipulated, or that there was only visual manipulation. This tendency may stem from visual primacy or cognitive load. Humans are highly visual creatures, and our brains prioritize visual stimuli over auditory signals. One potential factor may be sensory discrepancy; in some cases, deepfake videos exhibit more pronounced visual discrepancies than auditory ones. Furthermore, processing audio and visual information simultaneously may cause cognitive challenges; therefore, participants found it easier to focus on visual cues, encouraging them to prioritize visual information when detecting manipulations in videos. As shown in Fig. 9, of the 41% of “fake” responses, 16% believed that there were two manipulation modes, 15% believed that there was only visual manipulation, and only 10% believed that there was only auditory manipulation. However, audio manipulation accounts for 10% of all videos, and visual and audiovisual manipulation accounts for 20%. The results show that humans tend to underestimate the actual number of fake samples in detection.
@@ IV-B Detection Accuracy under Different Participant Groupings
Overall, t-test analysis showed that detection accuracy differed significantly between participants of different age groups and genders, while there were no significant differences in detection accuracy between different IT skill levels or between native and non-native English-speaking participants. Notably, participants showed improved performance in Phase II regardless of grouping, as shown in Fig. 10. The reasons have been discussed in the previous section.
@@
The study further analyzes the correlation between accuracy and confidence, as this insight is valuable for assessing participants’ metacognitive awareness and decision-making. To estimate whether participants correctly perceived their ability to detect audiovisual deepfakes, we compared their actual detection accuracy with self-reported confidence across all videos. Surprisingly, we found that participant’s self-reported confidence in their answers did not align with their accuracy in detecting audiovisual deepfake videos. Despite expressing high confidence, their actual accuracy was low. As shown in Fig. 12, most participants reported average confidence that was higher than their average accuracy. The average confidence level across all participants was 77.60%, while the average accuracy was 65.64%. This difference means that humans’ overconfidence in deepfake detection will cause deepfakes to bring greater harm to human society.
People tend to overestimate their ability to identify fake videos, which can lead to false assurances about the reliability of their judgments, poor decision-making, and increased susceptibility to misinformation. Several psychological and cognitive factors may contribute to this overestimation. People often fall victim to cognitive biases, such as the Dunning-Kruger effect [72], which often causes people to overestimate their skills in tasks for which they are not strong. Another factor may be that participants interpret their familiarity with certain faces or voices as evidence of authenticity [73]; this reliance may increase their confidence even if deepfakes manipulate cues they believe to be reliable. Likewise, some people believe that their exposure to the media or prior experience gives them control over discerning real from fake. This illusion of control [74] may lead to an overestimation of one’s own detection capabilities.
@@ IV-E Performance of AI Models
Model Class Precision Recall F1-Score Accuracy (%)
LipForensics [22] Real 0.95 0.90 0.92 92.5
Fake 0.90 0.95 0.93
AV-Lip-Sync [27] Real 0.94 0.80 0.87 87.5
Fake 0.83 0.95 0.88
AV-Lip-Sync+ [24] Real 1.00 0.95 0.97 97.5
Fake 0.95 1.00 0.98
CNN-Ensemble [25] Real 0.95 1.00 0.98 97.5
Fake 1.00 0.95 0.97
AVTENet [26] Real 0.95 1.00 0.98 97.5
Fake 1.00 0.95 0.97
@@
Fig. 14 compares the performance of humans and 5 AI models in detecting audiovisual deepfakes. Human performance was well below expectations and marginally above the chance level. However, all AI models performed significantly better than human participants. Arguably, AI models have demonstrated an impressive ability to detect subtle manipulations in audiovisual content that might escape human perception. We conducted two-sample t-tests to compare human performance against each of the five AI deepfake detection models. As shown in Table III, all AI models demonstrated significantly superior performance in audiovisual deepfake detection compared to humans, with p-values less than 0.05 and large negative t-values indicating a considerable performance gap. The results confirmed that AI models significantly outperformed humans. AI models excel humans at integrating information from multiple modalities, and combining audio and visual modalities in deepfake detection contributes significantly to the effectiveness of AI models.
TABLE III: Detection Accuracy of Human and AI models.
Method
Accuracy
t-value
p-value
Significance
$p<0.05$
Human 65.64 - - -
LipForensics [22] 92.50 -23.58 4.42 ×$\times 10^{-45}$ True
AV-Lip-Sync [27] 87.50 -19.54 6.40 ×$\times 10^{-38}$ True
AV-Lip-Sync+ [24] 97.50 -27.33 2.18 ×$\times 10^{-51}$ True
CNN-Ensemble [25] 97.50 -27.38 1.91 ×$\times 10^{-51}$ True
AVTENet [26] 97.50 -27.43 2.87 ×$\times 10^{-51}$ True
Comparisons between human observers and AI models highlight an important aspect of audiovisual deepfake detection. Despite the adaptability of human perception, AI models in this field consistently outperform humans. Humans were 65.64% accurate at detecting audiovisual deepfakes, indicating that humans have difficulty identifying forgeries in videos. In comparison, AI models’ accuracy was significantly higher, ranging from 87.5% to 97.5%. Possible key factors for AI models to perform well are as follows:
  • Consistency in Decision-Making: AI models follow predefined rules and decision thresholds to ensure absolute consistency in their detection strategies. However, humans suffer from fatigue, cognitive biases, and subjective inconsistencies that affect their decision-making.
  • Feature Extraction Capabilities: Humans rely on visual cues such as facial expressions, lip movements, and tone of voice in deepfake detection, but these features can be convincingly fake. However, AI models use advanced techniques such as CNNs, transformers, ensemble learning, or multimodal features of audio and video to detect subtle manipulations and flag deepfakes that look realistic to human observers.
  • Scalability and Learning from Data: AI models trained on FakeAVCeleb with variants of manipulated samples can effectively generalize or adapt to different fake generation techniques. In contrast, humans are hindered by limited exposure to deepfakes and cognitive biases. Especially with the rapid development of deepfake generation methods, without the help of automated tools, it is difficult for people’s judgment ability to keep up with the times.
  • Fatigue and Bias: Human detection capabilities tend to degrade over time due to fatigue and bias, especially when tasks are repeated. AI models ensure steady performance and no fatigue during large-scale content processing.
@@ VI Conclusions and Future Work
In this paper, we present the results of a subjective study aimed at analyzing human behavior in detecting audiovisual deepfake videos and comparing human performance with that of five AI models. We tested 110 human subjects and five AI models using the same set of 40 manually selected videos (20 real and 20 fake). The findings showed that subjects performed better than chance but generally found detecting deepfakes challenging. They performed marginally better than random chance, even when explicitly informed of the presence of deepfakes. The concern is that deepfake technology has reached a level of reality that could confuse much of the public, especially social media users, while people are overconfident in their ability to determine the authenticity of a video. However, over time, participants seemed to refine their perception and develop a more accurate understanding of their detection capabilities. Additionally, we found that familiarizing participants with the threats posed by deepfakes had no significant impact on their ability to detect deepfakes. Participants’ age, gender, and whether they were native speakers of English had an impact on their detection performance, but not their IT skill level. On the other hand, AI models outperformed human participants in detecting forgeries in videos. Furthermore, deepfake detection methods that integrate multiple modalities are more effective than unimodal methods. In order to reduce the harm of deepfakes to human society, it is crucial to develop deepfake detection technology.
This study has important implications for forensic analysis and adaptive countermeasures that improve the accuracy of digital forensic investigations, the attribution and tracking of manipulated content, the validation and verification of audiovisual evidence, and the development of forensic tools. Our findings highlight the need for improved training programs that incorporate AI support into forensic contexts; these insights can help law enforcement and media organizations develop more advanced detection tools and ensure better verification of audiovisual authenticity. Our findings also highlight the importance of adaptive countermeasures, such as real-time monitoring systems, that can dynamically keep pace with the evolution of deepfake technology to maintain the integrity of digital media in critical applications. This research offers a wide range of applications. In forensic analysis, these insights can drive the creation of better tools that combine human intuition with AI models. These tools will strengthen the fight against misinformation and will be valuable for media organizations, law enforcement, and legal professionals to verify audiovisual content.
To further expand and refine our understanding of human perception of audiovisual deepfakes, we can implement a test design that measures subjective performance in a more discreet and unbiased way. For example, the applicability of our findings can be generalized by including a broader and more diverse population. Detailed analysis of the auditory and visual components of audiovisual signals will reveal how individuals perceive, interpret and process audio and visual cues in multimodal deepfake videos. One potential area for future research is analyzing and understanding the cognitive biases that hinder the human ability to identify deepfakes. Furthermore, exploring the impact of the popularity of people in videos (e.g., celebrities or politicians) may provide valuable insights into enhancing human detection capabilities and understanding how prior knowledge affects the perception of corresponding deepfake content.